OpenStudy (anonymous):

A number is decreased by seven. Then, the new number is multiplied by 3 to get an answer of -9. What is the original number? -30 30 4 -4

OpenStudy (anonymous):

4

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

4-7 = -3*3=-9

OpenStudy (anonymous):

ohhh!! Thanks!! I will medal you!!

OpenStudy (anonymous):

no problem